Give the dimensions of rectangels with perimeters of 70 feet and length to width ratios of 3 to 4, 4 to 5 and 1 to 1

Answer by checkley77(12844)

2L+2W=P
-------
3N+4N=7N
7N=70
N=70/7
N=10
2L=30
2L=3N
2L=3*10
2L-30
L=30/2
L=15 FT. FOR THE LENGTH.
2W=4N
2W=4*10
2W=40
W=40/2
W=20 FT. FOR THE WIDTH.
PROOF:
2*15+2*20=70
30+40=70
-----------------------------
4N+5N=9N
9N=70
N=70/9
N=7.778
2L=4*7.778
2L=31.112
L=31.112/2
L=15.556 FT. FOR THE LENGTH.
2W=5*7.778
2W=38.89
W=38.89/2
W=19.445 FT FOR THE WIDTH.
PROOF:
2*15.556+2*19.445=70
31.112+38.89=70
70~70
---------------------------------
1N+1N=70
2N=70
N=70/2
N=35
2L=35
L=35/2
L=17.5 FT FOR THE LENGTH.
2W=45
W=35/2
W=17.5 FT. FOR THE WIDTH.
PROOF:
2*17.5+2*17.5=70
35+35=70
70=70